# Revealing the building blocks of tree balance: fundamental units of the Sackin and Colless Indices

### v1

@@ 3.1.2 Minimizing $\Delta_{CS}$ and maximizing $N_{b}$ as well as $\Delta_{SC}$
The main aim of this section is to give a full characterization of all trees that minimize $\Delta_{CS}$ and maximize both $\Delta_{SC}$ and $N_{b}$. And important step in this direction will be to show that every tree that minimizes the Colless index also minimizes $\Delta_{CS}=C-S$ (and thus also maximizes $N_{b}=-\frac{1}{2}\Delta_{CS}$). Note that this is highly non-obvious, because every Colless minimum also minimizes the Sackin index, cf. Section 2, so the term $S(T)$ that is subtracted from $C(T)$ also gets minimal in this case. However, the following recursion will be a crucial tool used to tackle the minimum in the following.
@@
Note that in $\mathcal{BT}^{*}_{0}$ and $\mathcal{BT}^{*}_{1}$ there is only one tree, respectively, namely $T^{\emptyset}$ and the tree consisting of only one vertex. In both cases, we have no inner vertices in the tree, implying that the sums defining $C(T)$ and $S(T)$ are empty, which shows that indeed we have $d_{0}=d_{1}=0-0=0$. To show the recursions $d_{2n}=2d_{n}-2n$ and $d_{2n+1}=d_{n+1}+d_{n}-2n$ for all $n\geq 1$, we perform a proof by induction on $n$. In the base case $n=1$, we need to calculate $d_{2\cdot 1}=d_{2}$ and $d_{2\cdot 1+1}=d_{3}$. However, note that there is only one tree each in $\mathcal{BT}^{*}_{2}$ and $\mathcal{BT}^{*}_{3}$ ($T^{cat}_{2}$ and $T^{cat}_{3}$, respectively), which come with Sackin indices $S(T^{cat}_{2})=2$ and $S(T^{cat}_{3})=5$ and Colless indices $C(T^{cat}_{2})=0$ and $C(T^{cat}_{3})=1$. Thus, we have:
$d_{2}=\Delta_{CS}(T^{cat}_{2})=C(T^{cat}_{2})-S(T^{cat}_{2})=0-2=-2=2\cdot 0-2=2\cdot d_{1}-2\cdot 1,$
@@
We proceed with the induction step, i.e., we assume that the recursions hold for all trees with $\tilde{n}<n$ leaves and now consider trees with $n$ leaves. We first show that $d_{2n}\leq 2d_{n}-2n$ and $d_{2n+1}\leq d_{n+1}+d_{n}-2n$ for all $n\in\mathbb{N}_{\geq 1}$. By Proposition 1 and Corollary 1, we have:
@@
To obtain equality in the recursions, it remains to show that $d_{2n}\geq 2d_{n}+2n$ and $d_{2n+1}\geq d_{n+1}+d_{n}+2n$ holds as well.
@@
$=\min\limits_{j\in\{1,\ldots,n\}}\left.\begin{cases}2d_{\frac{2n-j}{2}}-2\left(\frac{2n-j}{2}\right)+2d_{\frac{j}{2}}-2\frac{j}{2}-2j&\text{for $j$ even}\\ d_{\frac{2n-j+1}{2}}+d_{\frac{2n-j-1}{2}}-2\left(\frac{2n-j-1}{2}\right)+d_{\frac{j+1}{2}}+d_{\frac{j-1}{2}}-2\left(\frac{j-1}{2}\right)-2j&\text{for $j$ odd}\end{cases}\right\}$
$=\min\limits_{j\in\{1,\ldots,n\}}\left.\begin{cases}2d_{n-\frac{j}{2}}-2n+j+2d_{\frac{j}{2}}-j-2j&\text{for $j$ even}\\ d_{n-\frac{j-1}{2}}+d_{n-\frac{j+1}{2}}-2n+\underbrace{2\left(\frac{j+1}{2}\right)-2j}_{=-2(\frac{j+1}{2})+2}+d_{\frac{j+1}{2}}+d_{\frac{j-1}{2}}-2\left(\frac{j-1}{2}\right)&\text{for $j$ odd}\end{cases}\right\}$
@@
We want to show that the $A$, $B$ and $C$ are all at least as large as $d_{n}$, except for $C$ in case $j$ is odd and $j=n$, in which case we have $C\geq d_{n}-2$. This will show us that in all cases we have $2A-2n\geq 2d_{n}-2n$ and $B+C-2n+2\geq 2d_{n}-2n$. In order to see this, we first re-write the minimum defining $d_{n}$ as follows:
@@
    • We first analyze the case $j=1$. In this case, obviously $B=d_{n-0}+d_{0}-2\cdot 0=d_{n}+d_{0}=d_{n}$, and $C=d_{n-1}+d_{1}-2$. Thus, it is obvious that $B\geq d_{n}$, and we also have $C\geq d_{n}$ by Equation (4) (which can be seen using $i=1$). Note that this in particular implies that if $j=1$, we have $B+C-2n+2\geq d_{n}+d_{n}-2n+2=2d_{n}-2n+2>2d_{n}-2n$.
@@
      $C$ $=d_{n-\frac{n+1}{2}}+d_{\frac{n+1}{2}}-2\left(\frac{n+1}{2}\right)=d_{\frac{n-1}{2}}+d_{n-\frac{n-1}{2}}-2\left(\frac{n-1}{2}+1\right)$
@@
      Note that this in particular implies that if $j=n$ and $j$ is odd, we have $B+C-2n+2\geq d_{n}+(d_{n}-2)-2n+2=2d_{n}-2n$.
    • Lastly, if $j$ is odd and $3\leq j\leq n-1$, then $1\leq\left\lfloor\frac{j}{2}\right\rfloor=\frac{j-1}{2}<\frac{j+1}{2}=\left\lceil\frac{j}{2}\right\rceil\leq\left\lceil\frac{n-1}{2}\right\rceil\leq\left\lfloor\frac{n}{2}\right\rfloor$. Thus, we can use $i=\frac{j-1}{2}$ in Equation (4) to derive that $B\geq d_{n}$. Similarly, we can use $i=\frac{j+1}{2}$ in Equation (4) to derive that $C\geq d_{n}$. Note that this in particular implies that if $1<j<n$ and $j$ is odd, we have $B+C-2n+2\geq d_{n}+d_{n}-2n+2=2d_{n}-2n$.
@@
The following theorem is crucial as it leads to a family of $\Delta_{CS}$ minimal trees, namely all Colless minima. These trees are then automatically also $N_{b}$ and $\Delta_{SC}$ maximal. As a bonus, we can then easily derive the minimal value for $\Delta_{CS}$ and the maximal values for $N_{b}$ and $\Delta_{SC}$ from the minimal Colless and Sackin values.
@@
Let $n\in\mathbb{N}$ and $T\in\mathcal{BT}^{*}_{n}$. Moreover, let $\breve{s}(x)=\min_{z\in\mathbb{Z}}|x-z|$ be the triangle wave for $x\in\mathbb{R}$. Let $k_{n}=\lceil\log_{2}(n)\rceil$. Then, we have:
@@
  • $\Delta_{SC}(T)\geq S(T^{mb}_{n})-C(T^{mb}_{n})=-2^{k_{n}}+n\cdot(k_{n}+1)-\left(\sum\limits_{j=1}^{k_{n}-1}2^{j}\cdot\breve{s}(2^{-j}\cdot n)\right)$ and
@@
Proof.
We start with $\Delta_{CS}$, and our proof strategy is as follows: We first consider $\Delta_{CS}(T_{n}^{mb})$ and show that this term follows the same recursions as the ones given by Proposition 3. This proves that $\Delta_{CS}(T_{n}^{mb})=d_{n}$ for all $n\in\mathbb{N}$. As $T_{n}^{mb}$ is known to be Colless-minimal (cf. Section 2.2.4), this immediately shows that all Colless-minimal trees also achieve the minimal value $d_{n}$ of $\Delta_{CS}$, because obviously all Colless minima have the same Colless value. Also, it is known from the literature that all Colless minima are also Sackin minimal (cf. Section 2.2.4). As the minimal values both for Colless and Sackin are also already known (cf. Sections 2.2.3 and 2.2.4), this automatically gives us a value for $d_{n}=\Delta_{CS}(T_{n}^{mb})$.
Therefore, we start by considering $\Delta_{CS}(T_{n}^{mb})$ and prove that $\Delta_{CS}(T_{n}^{mb})=d_{n}$. For $n=0$ and $n=1$ we have $S(T_{n}^{mb})=C(T_{n}^{mb})=0$ and thus $\Delta_{CS}(T_{n}^{mb})=C(T_{n}^{mb})-S(T_{n}^{mb})=0-0=0=d_{n}$ for $n\in\{0,1\}$.
@@
This shows that indeed, $T_{n}^{mb}$, which is a known Colless minimal tree, also minimizes $\Delta_{CS}$ and thus, by Lemma 2, also maximizes $N_{b}$. As stated above, every Colless minimum is also a Sackin minimum, and the minimum values for Colless and Sackin are known (and identical for all trees assuming these values). Thus, we derive for every Colless minimal tree $T$ the term $d_{n}=\Delta_{CS}(T)=C(T)-S(T)=C(T_{n}^{mb})-S(T_{n}^{mb})$. Inserting the terms for the minimal values of the Sackin and Colless indices given by Sections 2.2.3 and 2.2.4 completes the statement for $\Delta_{CS}$.
@@
Hence, by Theorem 3, we know that all Colless minimal trees also minimize $\Delta_{CS}$ and thus maximize $N_{b}$ and $\Delta_{SC}$. However, these trees are not the only ones leading to these extrema. In order to derive a characterization of all $\Delta_{CS}$ minimal trees, an important step is given by the following proposition, which shows that the echelon tree $T_{n}^{be}$ also plays an important role for $\Delta_{CS}$, even though it is not generally an extremal tree neither for Sackin nor for Colless.
@@
For all $n\in\mathbb{N}$, the echelon tree $T_{n}^{be}$ minimizes $\Delta_{CS}$ and maximizes $N_{b}$ as well as $\Delta_{SC}$.
@@
Note that by definition of $k_{n}$, we have $n\in\{2^{k_{n}-1}+1,\ldots,2^{k_{n}}\}$. Based on this, we now distinguish two cases based on the two cases of the sizes of $n_{a}^{gfb}$ and $n_{b}^{gfb}$ as stated in Section 2.2.1.
@@
    $\ \ =\Delta_{CS}(T_{a}^{gfb})+\Delta_{CS}(T_{b}^{gfb})-2n_{b}^{gfb}$
[…]
    where we know the latter term to be minimal by Theorem 3. The result for $\Delta_{SC}$ then follows by using $\Delta_{SC}=-\Delta_{CS}$ and Theorem 3. Moreover, by Lemma 2, the equality of $\Delta_{CS}(T_{n}^{be})=\Delta_{CS}(T_{n}^{gfb})$ also gives us the desired result $N_{b}(T_{n}^{be})=N_{b}(T_{n}^{gfb})$, which is maximal again by Theorem 3. This completes the proof for the case $n\geq 3\cdot 2^{k_{n}-2}$.
@@
    If $n<3\cdot 2^{k_{n}-2}$, we know by Section 2.2.1 that $T_{b}^{gfb}=T_{k_{n}-2}^{fb}$, i.e., $n_{b}^{gfb}=2^{k_{n}-2}$ and $n_{a}=n-2^{k_{n}-2}$. However, $T_{a}^{be}=T_{k_{n}-1}^{fb}$ and $n_{b}^{be}=n-2^{k_{n}-1}$. $T_{n}^{gfb}$ and $T_{n}^{be}$ are schematically depicted in Figure 1, together with a third tree, namely $T^{\prime}$, which we consider here, too. Formally, $T^{\prime}=(T_{a}^{\prime},T_{b}^{\prime})$ is defined to be the tree with $T_{a}^{\prime}=\left(T_{n-2^{k_{n}-1}}^{be},T_{k_{n}-2}^{fb}\right)$ and $T^{\prime}_{b}=T_{k_{n}-2}^{fb}$.
@@
    We start by showing that $T_{a}^{\prime}$ is an echelon tree. Note that by definition, as $T_{n-2^{k_{n}-1}}^{be}$ is already an echelon tree, the entire subtree $T_{a}^{\prime}$ is echelon if and only if the $2^{k_{n}-2}$ leaves of the fully balanced tree $T_{k_{n}-2}^{fb}$ was the largest power of two fitting in $n-2^{k_{n}-1}+2^{k_{n}-2}=n-2^{k_{n}-2}$ many leaves – or, in other words, if the next power of two, namely $2^{k_{n}-1}$, is strictly larger than the number $n-2^{k_{n}-2}$ of leaves in subtree $T_{a}^{\prime}$.
    Suppose the latter statement is not true, i.e., suppose $n-2^{k_{n}-2}\geq 2^{k_{n}-1}$. This immediately implies $n\geq 2^{k_{n}-1}+2^{k_{n}-2}=3\cdot 2^{k_{n}-2}$, which contradicts the fact that we are in the case in which $n<3\cdot 2^{k_{n}-2}$. Hence, $2^{k_{n}-2}$ is the largest power of two that is smaller than or equal to $n-2^{k_{n}-2}$, showing that indeed $T_{a}^{\prime}$ is an echelon tree.
@@
    It remains to show that $\Delta_{CS}(T_{n}^{gfb})=\Delta_{CS}(T^{\prime})=\Delta_{CS}(T_{n}^{be})$, which will complete the contradiction for $\Delta_{CS}$ (and the required contradictions for $N_{b}$ and $\Delta_{SC}$ then automatically follow by applying $\Delta_{SC}=-\Delta_{CS}$, Lemma 2 and Theorem 3).
@@
    $\Delta_{CS}(T^{\prime})$ $=C(T^{\prime})-S(T^{\prime})$
    $=(C(T_{n}^{be})+n-2^{kn}+2^{k_{n}-2})-(S(T_{n}^{be})-2^{k_{n}}+n+2^{k_{n}-2})$
    $=C(T_{n}^{be})-S(T_{n}^{be})$
    $=\Delta_{CS}(T_{n}^{be}).$
    Hence, altogether we have $\Delta_{CS}(T_{n}^{gfb})=\Delta_{CS}(T^{\prime})=\Delta_{CS}(T_{n}^{be})$, a contradiction to our assumption that $\Delta_{CS}(T_{n}^{be})>\Delta_{CS}(T_{n}^{gfb})$. This completes the proof for $\Delta_{CS}$. The statements for $N_{b}$ and $\Delta_{SC}$ follow by Theorem 3, Lemma 2 and Remark 1.
@@
Let $T=(T_{a},T_{b})$ be a rooted binary tree with $n=n_{a}+n_{b}$ leaves with $n_{a}\geq n_{b}$ being the numbers of leaves in $T_{a}$ and $T_{b}$, respectively. Let $k_{n}=\lceil\log_{2}n\rceil$. Then, if $n_{a}>2^{k_{n}-1}$, $T$ is not $\Delta_{CS}$ minimal and not $N_{b}$ or $\Delta_{SC}$ maximal.
@@
Seeking a contradiction, assume there is a tree $T$ as described in the lemma which is indeed $\Delta_{CS}$ minimal (and thus, $\Delta_{SC}$ maximal, and – by Lemma 2 – also $N_{b}$ maximal). We consider the smallest $n$ for which such a tree $T$ exists.
As $T$ minimizes $\Delta_{CS}$ by assumption, by Corollary 1, $T_{a}$ and $T_{b}$ must minimize $\Delta_{CS}$, too. We now consider tree $\widetilde{T}$ as depicted in Figure 2, which is like $T$, but in which $T_{a}$ got exchanged by $T_{n_{a}}^{be}$.
@@
$S(T^{\prime})$ $=S(\widetilde{T})-n_{a}+(n_{a}-2^{k_{n}-1}+n_{b})$
$=S(\widetilde{T})-2^{k_{n}-1}+n_{b},$
@@
$=\Delta_{CS}(\widetilde{T})-n_{a}-n_{b}+2^{k_{n}-1}+|n_{a}-2^{k_{n}-1}-n_{b}|.$ (4)
We now distinguish two cases: It may be that $T^{\ast}$ contains at least as many leaves as $T_{b}$, in which case we have $n_{a}-2^{k_{n}-1}\geq n_{b}$ and thus $|n_{a}-2^{k_{n}-1}-n_{b}|=n_{a}-2^{k_{n}-1}-n_{b}$. Or, it might be that $n_{a}-2^{k_{n}-1}<n_{b}$ and thus $|n_{a}-2^{k_{n}-1}-n_{b}|=2^{k_{n}-1}+n_{b}-n_{a}$. In the first case, using Equation (4), we get:
[…]
In the second case, however, Equation (4) leads to:
@@
Before we can finally show that $\Delta_{CS}$ is an imbalance index and $N_{b}$ and $\Delta_{SC}$ are balance indices, we still need to prove the following proposition.

### v2

@@ 3.1.2 Minimizing $\Delta_{CS}$ and maximizing $N_{b}$ as well as $\Delta_{SC}$
The main aim of this section is to give a full characterization of all trees that minimize $\Delta_{CS}$. An important step in this direction will be to show that every tree that minimizes the Colless index also minimizes $\Delta_{CS}=C-S$. Note that this is highly non-obvious, because every Colless minimum also minimizes the Sackin index, cf. Section 2, so the term $S(T)$ that is subtracted from $C(T)$ also gets minimal in this case. However, the following recurrence will be an essential tool used to tackle the minimum in the following.
@@
Before proving the recurrences for $d_{2n}$ and $d_{2n+1}$ for $n\in\mathbb{N}_{\geq 1}$, we briefly show that the initial values $d_{0}=d_{1}=0$ are correct. Note that in $\mathcal{BT}^{*}_{0}$ and $\mathcal{BT}^{*}_{1}$ there is only one tree, respectively, namely $T^{\emptyset}$ and the tree consisting of only one vertex. In both cases, we have no inner vertices in the tree, implying that the sums defining $C(T)$ and $S(T)$ are empty, which shows that indeed we have $d_{0}=d_{1}=0-0=0$.
We now turn our attention to the recurrences. To show that $d_{2n}=2d_{n}-2n$ and $d_{2n+1}=d_{n+1}+d_{n}-2n$ for all $n\geq 1$, we perform a proof by induction on $n$. In the base case $n=1$, we need to calculate $d_{2\cdot 1}=d_{2}$ and $d_{2\cdot 1+1}=d_{3}$. However, note that there is only one tree each in $\mathcal{BT}^{*}_{2}$ and $\mathcal{BT}^{*}_{3}$ ($T^{cat}_{2}$ and $T^{cat}_{3}$, respectively), which come with Sackin indices $S(T^{cat}_{2})=2$ and $S(T^{cat}_{3})=5$ and Colless indices $C(T^{cat}_{2})=0$ and $C(T^{cat}_{3})=1$. Thus, we have:
$d_{2}=\Delta_{CS}(T^{cat}_{2})=C(T^{cat}_{2})-S(T^{cat}_{2})=0-2=2\cdot d_{1}-2\cdot 1,$
@@
We proceed with the induction step, i.e., we assume that the recurrences hold for all trees with $\tilde{n}<n$ leaves and now consider trees with $n$ leaves. We first show that $d_{2n}\leq 2d_{n}-2n$ and $d_{2n+1}\leq d_{n+1}+d_{n}-2n$ for all $n\in\mathbb{N}_{\geq 1}$. By Proposition 1 and Corollary 1, we have:
@@
To obtain equality in the recurrences, it remains to show that $d_{2n}\geq 2d_{n}-2n$ and $d_{2n+1}\geq d_{n+1}+d_{n}-2n$ holds as well.
@@
We want to show that $A$, $B$ and $C$ are all at least as large as $d_{n}$, except for $C$ in case $j$ is odd and $j=n$, in which case we have $C\geq d_{n}-2$. This will show that in all cases we have $2A-2n\geq 2d_{n}-2n$ and $B+C-2n+2\geq 2d_{n}-2n$. In order to see this, we first re-write the minimum defining $d_{n}$ as follows:
@@
    • We first analyze the case $j=1$. In this case, obviously $B=d_{n}$, and $C=d_{n-1}+d_{1}-2$. Thus, it is obvious that $C\geq d_{n}$ by Equation (4) (which can be seen using $i=1$).
@@
    • Lastly, if $j$ is odd and $3\leq j\leq n-1$, then $1\leq\left\lfloor\frac{j}{2}\right\rfloor=\frac{j-1}{2}<\frac{j+1}{2}=\left\lceil\frac{j}{2}\right\rceil\leq\left\lceil\frac{n-1}{2}\right\rceil\leq\left\lfloor\frac{n}{2}\right\rfloor$. Thus, we can use $i=\frac{j-1}{2}$ in Equation (4) to derive that $B\geq d_{n}$. Similarly, we can use $i=\frac{j+1}{2}$ in Equation (4) to derive that $C\geq d_{n}$.
    Hence, in case $j$ is odd and $1\leq j\leq n$, we have $B+C-2n+2\geq 2d_{n}-2n$.
@@
The following theorem provides a family of $\Delta_{CS}$ minimal trees, namely all Colless minima. As a bonus, we can then easily derive the minimal value for $\Delta_{CS}$ and the maximal values for $N_{b}$ and $\Delta_{SC}$ from the minimal Colless and Sackin values.
@@
Let $n\in\mathbb{N}$ and $T\in\mathcal{BT}^{*}_{n}$. Moreover, let $\breve{s}(x)=\min_{z\in\mathbb{Z}}|x-z|$ for $x\in\mathbb{R}$. Let $k_{n}=\lceil\log_{2}(n)\rceil$. Then, we have:
@@
  • $\Delta_{SC}(T)\leq S(T^{mb}_{n})-C(T^{mb}_{n})=-2^{k_{n}}+n\cdot(k_{n}+1)-\left(\sum\limits_{j=1}^{k_{n}-1}2^{j}\cdot\breve{s}(2^{-j}\cdot n)\right)$ and
@@
Our proof strategy for $\Delta_{CS}$ will be as follows: We first consider $\Delta_{CS}(T_{n}^{mb})$ and show that this term follows the same recurrences as the ones given by Proposition 3. This proves that $\Delta_{CS}(T_{n}^{mb})=d_{n}$ for all $n\in\mathbb{N}$. As $T_{n}^{mb}$ is known to be Colless-minimal (cf. Section 2.2.4), this immediately shows that all Colless-minimal trees also achieve the minimal value $d_{n}$ of $\Delta_{CS}$, because obviously all Colless minima have the same Colless value. Also, it is known from the literature that all Colless minima are also Sackin minima (cf. Section 2.2.4). As the minimal values both for Colless and Sackin are also already known (cf. Sections 2.2.3 and 2.2.4), this automatically gives us a value for $d_{n}=\Delta_{CS}(T_{n}^{mb})$.
Proof of Theorem 3.
We start by considering $\Delta_{CS}(T_{n}^{mb})$ and prove that $\Delta_{CS}(T_{n}^{mb})=d_{n}$. For $n=0$ and $n=1$ we have $S(T_{n}^{mb})=C(T_{n}^{mb})=0$ and thus $\Delta_{CS}(T_{n}^{mb})=C(T_{n}^{mb})-S(T_{n}^{mb})=0-0=0=d_{n}$ for $n\in\{0,1\}$.
@@
This shows that indeed, $T_{n}^{mb}$, which is a known Colless minimal tree, also minimizes $\Delta_{CS}$. As stated above, every Colless minimum is also a Sackin minimum, and the minimum values for Colless and Sackin are known (and identical for all trees assuming these values). Thus, we derive for every Colless minimal tree $T$ the term $d_{n}=\Delta_{CS}(T)=C(T)-S(T)=C(T_{n}^{mb})-S(T_{n}^{mb})$. Inserting the terms for the minimal values of the Sackin and Colless indices given by Sections 2.2.3 and 2.2.4 completes the statement for $\Delta_{CS}$.
@@
Hence, by Theorem 3, we know that all Colless minimal trees also minimize $\Delta_{CS}$. However, these trees are not the only ones leading to these extrema. In order to derive a characterization of all $\Delta_{CS}$ minimal trees, an important step is given by the following proposition, which shows that the echelon tree $T_{n}^{be}$ also plays another important role for $\Delta_{CS}$, even though it is not generally an extremal tree neither for Sackin nor for Colless.
@@
For all $n\in\mathbb{N}$, the echelon tree $T_{n}^{be}$ minimizes $\Delta_{CS}$ (and maximizes $N_{b}$ and $\Delta_{SC}$).
@@
Note that by definition of $k_{n}$, we have $n\in\{2^{k_{n}-1}+1,\ldots,2^{k_{n}}\}$. Based on this, we now distinguish two cases based on the two possible sizes of $n_{a}^{gfb}$ and $n_{b}^{gfb}$ as stated in Section 2.2.1.
@@
    where we know the latter term to be minimal by Theorem 3. This completes the proof for the case $n\geq 3\cdot 2^{k_{n}-2}$.
@@
    If $n<3\cdot 2^{k_{n}-2}$, we know by Section 2.2.1 that $T_{b}^{gfb}=T_{k_{n}-2}^{fb}$, i.e., $n_{b}^{gfb}=2^{k_{n}-2}$ and $n_{a}=n-2^{k_{n}-2}$. However, $T_{a}^{be}=T_{k_{n}-1}^{fb}$ and $n_{b}^{be}=n-2^{k_{n}-1}$. $T_{n}^{gfb}$ and $T_{n}^{be}$ are schematically depicted in Figure 1, together with a third tree, namely $T^{\prime}$, which we consider here, too. Formally, $T^{\prime}=(T_{a}^{\prime},T_{b}^{\prime})$ is defined to be the tree with $T_{a}^{\prime}=\left(T_{n-2^{k_{n}-1}}^{be},T_{k_{n}-2}^{fb}\right)$ and $T^{\prime}_{b}=T_{k_{n}-2}^{fb}$. Note that $T^{\prime}$ is such that its root implies the gfb partition and one of its maximum pending subtrees coincides with the gfb tree’s smaller one. Additionally, its larger maximum pending subtree’s two maximum pending subtrees together with this smaller subtree are identical to the three subtrees employed by the echelon tree, cf. Figure 1. Thus, $T^{\prime}$ can be regarded as a mixture of $T_{n}^{gfb}$ and $T_{n}^{be}$.
@@
    We start by showing that $T_{a}^{\prime}$ is an echelon tree. Note that $n-2^{k_{n}-1}+2^{k_{n}-2}=n-2^{k_{n}-2}<2^{k_{n}-1}$ as $n<3\cdot 2^{k_{n}-2}$ by assumption. This shows that $2^{k_{n}-2}$ is the largest power of two fitting into the $n-2^{k_{n}-2}$ leaves of $T_{a}^{\prime}$, and with $T_{a}^{\prime}=\left(T_{k_{n}-2}^{fb},T_{n-2^{k_{n}-1}}^{be}\right)$ it is clear that $T_{a}^{\prime}$ is indeed an echelon tree.
@@
    It remains to show that $\Delta_{CS}(T_{n}^{gfb})=\Delta_{CS}(T^{\prime})=\Delta_{CS}(T_{n}^{be})$, which will complete the contradiction for $\Delta_{CS}$.
@@
    $\Delta_{CS}(T^{\prime})$ $=C(T^{\prime})-S(T^{\prime})=\underbrace{(C(T_{n}^{be})+n-2^{kn}+2^{k_{n}-2})}_{=C(T_{n}^{be})}-\underbrace{(S(T_{n}^{be})-2^{k_{n}}+n+2^{k_{n}-2})}_{=S(T_{n}^{be})}=\Delta_{CS}(T_{n}^{be}).$
    Hence, altogether we have $\Delta_{CS}(T_{n}^{gfb})=\Delta_{CS}(T^{\prime})=\Delta_{CS}(T_{n}^{be})$, a contradiction to our assumption that $\Delta_{CS}(T_{n}^{be})>\Delta_{CS}(T_{n}^{gfb})$. This completes the proof.
@@
Let $T=(T_{a},T_{b})$ be a rooted binary tree with $n=n_{a}+n_{b}$ leaves with $n_{a}\geq n_{b}$ being the numbers of leaves in $T_{a}$ and $T_{b}$, respectively. Let $k_{n}=\lceil\log_{2}n\rceil$. Then, if $n_{a}>2^{k_{n}-1}$, $T$ is not $\Delta_{CS}$ minimal.
@@
Seeking a contradiction, assume there is a tree $T$ as described in the lemma which is indeed $\Delta_{CS}$ minimal. We consider the smallest $n$ for which such a tree $T$ exists. As $T$ minimizes $\Delta_{CS}$ by assumption, by Corollary 1, $T_{a}$ and $T_{b}$ must minimize $\Delta_{CS}$, too. We now consider tree $\widetilde{T}$ as depicted in Figure 2, which is like $T$, but in which $T_{a}$ got exchanged by $T_{n_{a}}^{be}$.
@@
$S(T^{\prime})$ $=S(\widetilde{T})-n_{a}+(n_{a}-2^{k_{n}-1}+n_{b})=S(\widetilde{T})-2^{k_{n}-1}+n_{b},$
@@
$=\Delta_{CS}(\widetilde{T})-(n_{a}+n_{b}-2^{k_{n}-1})+\max\{\underbrace{n_{a}-2^{k_{n}-1}-n_{b}}_{<n_{a}+n_{b}-2^{k_{n}-1}},\underbrace{-n_{a}+2^{k_{n}-1}+n_{b}}_{\begin{subarray}{c}<n_{a}+n_{b}-2^{k_{n}-1}\\ \mbox{ \small as }n_{a}>2^{k_{n}-1}\end{subarray}}\}$
@@
Before we can finally show that $\Delta_{CS}$ is an imbalance index, we still need to prove the following proposition.